### v2

@@ Abstract
Advances in peptide identification are revolutionizing our ability to decipher protein functions and accelerate therapeutic discovery. We present PDeepPP, a deep learning framework that integrates pretrained protein language models with parallel transformer-CNN architectures, achieving state-of-the-art performance in peptide characterization tasks. The model’s hybrid architecture demonstrates unique capabilities in capturing both local sequence motifs and global structural features, as evidenced by 29% improved cluster separation in UMAP visualizations compared to conventional approaches. Evaluated across 33 biological recognition tasks - including post-translational modification site prediction and bioactive peptide identification - PDeepPP outperformed existing methods in 25 tasks with average AUC improvements of 4.2%. Notably, it achieved 0.9726 accuracy with PR AUC 0.9977 in antimicrobial peptide detection while reducing false negatives by 37.5% in antimalarial recognition scenarios. This framework enables accurate large-scale peptide analysis, achieving 218× acceleration over sequence-alignment-based methods while maintaining 99.5% specificity in critical glycosylation site detection.PDeepPP establishes a new paradigm for computational peptide analysis through its synergistic architecture design, enabling rapid yet precise functional annotation that bridges molecular pattern recognition with translational biomedical applications.We have made our implementation, including code, data, and pretrained models, publicly available via GitHub (https://github.com/fondress/PDeepPP) and Hugging Face (https://huggingface.co/fondress/PDeppPP).
@@ 1 Introduction
The identification of functional peptides is crucial in computational biology due to their significant roles in therapeutic and industrial applicationsdeep-stp ; malignant . Unlike traditional experimental methods, which are labor-intensive, costly, and time-consuming, computational approaches provide a more efficient alternative. However, existing deep learning-based methods for peptide identification often suffer from excessive complexity, inconsistency, and limited generalizability across diverse datasetsApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer . Addressing these challenges is essential for advancing drug development, functional annotation, and biomarker discovery.
Among the key areas in functional peptide identification are protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 . PTMs, including phosphorylationphosphorylation , glycosylationGlycosylation , and acetylationAcetylation , play vital roles in regulating protein function and cellular signaling, influencing various health and disease processesPTMindisease . Similarly, BPs, such as antimicrobial, anticancer, and antihypertensive peptides, hold significant potential in pharmaceutical applicationsBioactivePeptides ; BP5 ; BP6 . However, the experimental identification of these peptides is hindered by high costs, extensive labor, and prolonged time requirementschallengeofPTM ; zhipu ; zhipucuowu ; ProcessingBioactivePeptide ; food_protein-derivedBPs .
To overcome these limitations, deep learning-based computational methods have been developed, including MusiteDeepMusite ; musite2 ; musite3 ; musite4 , UniDL4BioPepUnidl4Biopep , and various modification-specific predictorsPtrainsips ; Succinylation ; methy . Nevertheless, these approaches face challenges related to predictive performance, dataset scalability, and the lack of a unified frameworkprotein——DL ; ProgressforPTMprediction .
In this study, we propose PDeepPP, a novel deep learning framework that unifies the task of functional peptide identification across 37 tasks (33 benchmark datasets plus 4 ablation study datasets). PDeepPP integrates pretrained protein language models (e.g., ESM-2esm ) with a combination of transformertransformer and convolutional neural network (CNN)CNN architectures, enhanced by a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 to effectively handle imbalanced datasets. This unified approach not only simplifies the model architecture but also addresses the inconsistencies in existing methods.
@@
Furthermore, PDeepPP achieves 92.4% average recall under data-scarce conditions ($n<500$) through transfer learning, outperforming traditional methods by 34.7%. Its parallelized architecture processes population-scale proteomic data at a speed of 15,000 sequences per minute, offering a 218× acceleration compared to sequence-alignment-based approaches. This efficiency makes PDeepPP a powerful tool for large-scale functional annotation and targeted therapeutic discovery.
In summary, PDeepPP addresses the key challenges in functional peptide identification by providing a unified, efficient, and scalable solution, paving the way for the next generation of intelligent protein analysis.
@@ 2.1 Benchmark datasets
All benchmark datasets are sourced from existing review papers and datasets used for single-task predictions. We conducted a fair evaluation of metrics on the collected datasets. In total, we collected 37 datasets from two review articles and four single-task papers,including twenty BPs datasets for seventeen different bioactivities and 17 PTMs datasets for different modifications.Specificly contains angiotensinconverting enzyme (ACE) inhibitory activity (anti-hypertension)ACE , dipeptidyl peptidase (DPPIV) inhibitory activity (anti-diabetes)DPPIV , bitter bitter , umamiumami , antimicrobial activity antimicrobial , antimalarial activityantimalarial , quorum-sensing (QS) activityquorum.bibtex , anticancer activityanticancer1 ; anticancer2 , anti-methicillin-resistant S. aureus (MRSA)strains activity antimrsa , tumor T cell antigens (TTCA)TTCA , blood–brain barrierBBP , antiparasitic activity antiparastic , neuropeptidenuero1 ; nuero2 , antibacterial activity Antibacterial,Antifungal,Antiviral , antifungal activity Antibacterial,Antifungal,Antiviral , antiviralactivityAntibacterial,Antifungal,Antiviral , toxicitytoxicity and antioxidantantioxidant activity ;The following data are all derived from the UniProtKB/Swiss-Prot databaseuniprot : Phosphoserine/threonine, Phosphotyrosine, N-linked glycosylation, O-lined glycosylation, N6-acetyllysine, Methyllysine,
S-palmitoylation-cysteine, Pyrrolidone-carboxylic-acid,
Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R ,Ubiquitin_Kubiquitin and histone lysine crotonylationkcr . SuppTable 1 summarizes the basic information about the sources of the datasets. More information and the complete data can be found on [github].
@@ 2.2 Structure overview
Recent hybrid architectures hybrid ; hybrid2 demonstrate that combining ESM with task-specific embeddings improves PTM generalization. PDeepPP uses a parallel neural network with CNNs and Tranformers for module combination. The TransLinear layer uses the combination of encoder and fully connected layer, and the PosCNN layer uses the combination of position encoding and CNN. The outputs of the two networks are concatenated, followed by two convolutions to provide the predicted result. In terms of pre-training, PDeepPP uses the latest esm-2 for proteins to perform weighted combination with the training model and base embedding based on the fully connected layer to obtain better feature representation.The entire process used by PDeepPP is referenced in Fig. 1.
@@ 2.2.1 Data processing
For the benchmark data set, the original data set that was not divided was taken out according to the ratio of 20%. New results show that long sequences of proteins can enhance learning of molecular interactions,longtext so the sequence was cut at the same time, and every 33 consecutive amino acids were cut into a sequence.Sequences of insufficient length are padded with X, and partial overlap of sequences is allowed. For PTM sites, the positive site should be placed in the middle of the sequence. If there is a peptide chain of less than 33 amino acids, its ends are padded to the target length before training. The validation set is randomly selected at training time according to 10% of the training set.
@@ 2.2.2 Embedding strategy
The model used in this work is ESM-2 with 650 million parameters, a large-scale pre-trained model designed to capture the complex relationships within protein sequences through a Transformer architecture. ESM-2 is pre-trained using a Masked Language Modeling (MLM) objective on a large protein sequence dataset, enabling it to learn both local and global sequence dependencies. The model generates token-wise embeddings that provide rich contextual representations of proteins, which are particularly effective for tasks such as protein structure prediction and functional annotation.
In this study, we combine the pre-trained ESM-2 embeddings with a custom BaseEmbedding model to create a hybrid representation of protein sequences.BaseEmbedding pre-training is an embedding layer with a separate teleprompter that generates additional sequence representations, which are linearly transformed to match the ESM-2 output dimension (1280). This combination aims to leverage the pre-trained knowledge from ESM-2 while allowing the model to learn task-specific embeddings through the BaseEmbedding.
Moreover, residual connections are introduced between these two representations to facilitate feature refinement and ensure efficient gradient propagation, a design choice shown to improve training stability and overall performance in deep neural networksresnet1 ; resnet2 .
The combination of these two representations is controlled by a predefined ratio (esmratio), which is set to 0.9 in this case. This means that 90% of the final representation comes from the ESM-2 embeddings, while 10% comes from the BaseEmbedding model. The final combined sequence representation is computed using the following weighted sum:
@@
$\text{R}_{\text{combined}}$ represents The combined sequence feature representation; $\alpha$ (set to 0.9) is the ‘esm_ratio’, which defines the contribution of the ESM-2 embeddings; $\text{R}_{\text{ESM-2}}$ is the embedding generated by the ESM-2 model; $\text{R}_{\text{Base}}$ is the embedding generated by the BaseEmbedding model.
@@
Protein sequences are passed through the pre-trained ESM-2 model, which generates token-wise embeddings from its final layer. ESM-2 650M, based on the Transformer architecture, contains 32 layers with a hidden dimension of 1024. The model uses 16 attention heads, each of which operates with a dimensionality of 64. The feedforward network has a dimension of 2048. The model’s input layer embeds protein sequences into 1024-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence. These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function prediction. The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context. The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein identification, function annotation, and structure prediction.preexplanation
BaseEmbedding Model
The BaseEmbedding model consists of a learnable embedding layer that maps each amino acid in the sequence to a 128-dimensional high-dimensional vector. This vector is subsequently transformed through a linear layer to produce a 1280-dimensional representation, aligning it with the dimensionality of the ESM-2 embeddings.
@@
The embeddings generated by the ESM-2 model and the BaseEmbedding model are combined using the predefined ratio (esm_ratio = 0.9). The combined sequence representation is computed as shown in the formula above, where the ESM-2 embeddings contribute 90% of the final representation and the BaseEmbedding model contributes 10%.
Saving Representations
The combined representations for both the training and test datasets are saved as .npy files for use in downstream tasks. The corresponding labels are also saved, ensuring that the representations can be easily loaded and utilized for further model training or evaluation. This approach, which combines a powerful pre-trained model with a learnable embedding layer, provides a flexible and effective method for representing protein sequences. It allows the model to adapt to specific tasks while benefiting from the pre-trained knowledge encoded in the ESM-2 model.
@@ 2.2.3 TransConv1d Network
The TransConv1d layer combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data, making it suitable for processing sequential data or high-dimensional features. Its architecture consists of multiple modules, each with specific hyperparameters to ensure the model’s expressive power and training stability.
First, the input data is passed through the SelfAttention
-GlobalFeatures module, which uses multi-head self-attention (MultiheadAttention) to capture global features of the input data. Specifically, the [embed_dim] is set to the input feature dimension, and [num_heads] is set to 8 to parallelize the processing of different feature patterns. The output of the self-attention is added to the original input via a residual connection and then stabilized with layer normalization (LayerNorm). The output is then processed through two fully connected layers (fc1 and fc2), where the first layer maps the feature dimension to 256, and the second layer maps it to the final output size. Dropout is applied after the first fully connected layer to prevent overfitting.
By combining self-attention mechanisms, the Transformer encoder, and fully connected layers, the TransConv1d layer effectively captures the complex dependencies in the input data. The residual connections and layer normalization help maintain data stability and preserve information, thus improving the model’s expressive power and training stability.
@@ 2.2.4 PosCNN Network
The PosCNN layer integrates convolutional neural network (CNN) operations with optional positional encoding, designed to extract local features from input data while preserving sequence information, making it suitable for processing sequential data or high-dimensional features. The architecture focuses on capturing local patterns through convolution operations, followed by a fully connected layer for feature mapping.
The overall structure of the PosCNN layer is designed to extract local features from the input sequence, enhance these features with positional information (if enabled), and finally map them to the output space using a fully connected layer.
@@ 2.2.5 PredictModule
The forward process of the PredictModule first processes the input data x through the TransConv1d and PosCNN layers to obtain two feature representations. The outputs are then adjusted in dimension and concatenated along the feature dimension, forming a unified representation that combines both global and local features. The concatenated features are then permuted to match the input format required by the subsequent convolutional layers. The combined output from TransConv1d and PosCNN is passed through a convolutional layer with 32 output channels, followed by adaptive max pooling to reduce the sequence length to 1, with dropout applied for regularization. Next, the processed features are passed through another convolutional layer with 64 output channels, followed by max pooling and dropout. Finally, the features are flattened and input into a fully connected layer to generate the final prediction output. This sequence of operations helps refine the features, prevent overfitting, and produce robust prediction results.

### v3

@@ Abstract
Accurate identification of bioactive peptides (BPs) and protein post-translational modifications (PTMs) is essential for understanding protein function and advancing therapeutic discovery. However, most computational methods remain limited in their generalizability across diverse peptide functions. Here, we present PDeepPP, a unified deep learning framework that integrates pretrained protein language models with a hybrid transformer-convolutional architecture, enabling robust identification across diverse peptide classes and PTM sites. We curated comprehensive benchmark datasets and implemented strategies to address data imbalance, allowing PDeepPP to systematically extract both global and local sequence features. Through extensive analyses—including dimensionality reduction and comparison studies—PDeepPP demonstrates strong, interpretable peptide representations and achieves state-of-the-art performance in 25 of the 33 biological identification tasks. Notably, PDeepPP attains high accuracy in antimicrobial (0.9726) and phosphorylation site (0.9984) identification, with 99.5% specificity in glycosylation site prediction and substantial reduction in false negatives in antimalarial tasks. By enabling large-scale, accurate peptide analysis, PDeepPP supports biomedical research and the discovery of novel therapeutic targets for disease treatment. All code, datasets, and pretrained models are publicly available via GitHub (https://github.com/fondress/PDeepPP) and Hugging Face (https://huggingface.co/fondress/PDeppPP).
@@ 1 Introduction
The identification of functional peptides is crucial in computational biology due to their significant roles in therapeutic and industrial applicationsdeep-stp ; malignant . In contrast with traditional experimental methods, which are labor-intensive, costly, and time-consuming, computational approaches provide a more efficient alternative. However, existing deep learning-based methods for peptide identification often suffer from excessive complexity, inconsistency, and limited generalizability across diverse datasetsApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer . Addressing these challenges is essential for advancing drug development, functional annotation, and biomarker discovery.
Among the key areas in functional peptide identification are bioactive peptides (BPs)BP1 ; BP2 and protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 . Bioactive peptides, such as antimicrobial, anticancer, and antihypertensive peptides, hold significant potential in pharmaceutical applicationsBioactivePeptides ; BP5 ; BP6 . PTMs—including phosphorylationphosphorylation , glycosylationGlycosylation , and acetylationAcetylation —play vital roles in regulating protein function and cellular signaling, influencing various health and disease processesPTMindisease . However, the experimental identification of these peptides is hindered by high costs, extensive labor, and prolonged time requirementschallengeofPTM ; zhipu ; zhipucuowu ; ProcessingBioactivePeptide ; food_protein-derivedBPs .
To overcome these limitations, deep learning-based computational methods have been developed, including UniDL4BioPepUnidl4Biopep , MusiteDeepMusite ; musite2 ; musite3 ; musite4 , PhosF3Cphosf3c , and various modification-specific predictorsPtrainsips ; Succinylation ; methy . Nevertheless, these approaches face challenges related to predictive performance, dataset scalability, and the lack of a unified frameworkprotein——DL ; ProgressforPTMprediction . For example, while PhosF3C shows promise in phosphorylation site prediction, its performance on complex PTM patterns remains limited, with an average accuracy of 0.9514 across different PTM types.
In this study, we propose PDeepPP, a novel deep learning framework that unifies the task of functional peptide identification across 37 tasks (33 benchmark datasets plus 4 ablation study datasets). To improve robustness and generalizability, PDeepPP integrates pretrained protein language models (e.g., ESM-2esm ) and leverages transfer learning to extract rich contextual features from large-scale protein sequences. The framework combines transformertransformer and convolutional neural network (CNN)CNN architectures, further enhanced by a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 for effectively handling imbalanced datasets. This unified approach not only simplifies the model architecture but also addresses the inconsistencies in existing methods.
@@
PDeepPP addresses key challenges in functional peptide identification by providing a unified and scalable solution. In the control of the loss function combined with transfer learning, it achieves 92.4% average recall under data-scarce conditions ($n<500$), outperforming traditional methods by 34.7% in recall metrics. For instance, in antimicrobial peptide datasets with fewer than 500 positive samples, PDeepPP achieves a recall of 91.7%, while UniDL4BioPep and MusiteDeep attain 62.3% and 58.9%, respectively. In N6-acetyllysine modification prediction, the model reduces false negatives by 41% compared to PhosF3C, demonstrating its robustness in low-data scenarios. These advances make PDeepPP a promising tool for large-scale functional annotation—enabling 89.3% average accuracy across 17 PTM datasets with imbalanced ratios (up to 1:30 negative/positive samples)—and targeted therapeutic discovery, paving the way for the next generation of intelligent protein analysis.
@@ 2.1 Benchmark datasets
We compiled a comprehensive benchmark of 37 peptide prediction datasets, drawn from two review studies and four specialized publications. These include 20 datasets for various bioactivities and 17 datasets for distinct post-translational modifications. All selected datasets come from curated sources (e.g., Swiss-Prot-derived data), ensuring high quality and representativeness across diverse tasks. Specifically contains angiotensin converting enzyme (ACE) inhibitory activity (anti-hypertension)ACE , dipeptidyl peptidase (DPPIV) inhibitory activity (anti-diabetes)DPPIV , bitter bitter , umamiumami , antimicrobial activityantimicrobial , antimalarial activityantimalarial , quorum-sensing (QS) activityquorum.bibtex , anticancer activityanticancer1 ; anticancer2 , anti-methicillin-resistant S. aureus (MRSA) strains activity antimrsa , tumor T cell antigens (TTCA)TTCA , blood–brain barrierBBP , antiparasitic activity antiparastic , neuropeptidenuero1 ; nuero2 , antibacterial activity Antibacterial , antifungal activity Antibacterial , antiviral activityAntibacterial , toxicitytoxicity and antioxidantantioxidant activity. The following data are all derived from the UniProtKB/Swiss-Prot databaseuniprot : Phosphoserine/threonine, Phosphotyrosine, N-linked glycosylation, O-linked glycosylation, N6-acetyllysine, Methyllysine, S-palmitoylation-cysteine, Pyrrolidone-carboxylic-acid, Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R , Ubiquitin_Kubiquitin and histone lysine crotonylationkcr .
Detailed information about all benchmark datasets, including sample sizes, data sources, and specific task descriptions, is provided in Supplementary Material (SuppTable 1). The complete datasets are available on our GitHub repository.
@@ 2.2 Structure overview
Recent hybrid architectures hybrid ; hybrid2 demonstrate that combining ESM with task-specific embeddings improves PTM generalization. PDeepPP uses a parallel neural network with CNNs and Tranformers for module combination. The TransLinear module uses the combination of encoder and fully connected module, and the PosCNN module uses the combination of position encoding and CNN. The outputs of the two networks are concatenated, followed by two convolutions to provide the predicted result. In terms of pre-training, PDeepPP uses the latest esm-2 for proteins to perform weighted combination with the training model and base embedding based on the fully connected layer to obtain better feature representation.The entire process adopted by PDeepPP is referenced in Fig. 1.
@@ 2.2.1 Data processing
For the benchmark datasets, 20% of the original unsplit data was allocated as the test set. New results show that long sequences of proteins can enhance learning of molecular interactions,longtext so the sequence was cut at the same time, and every 33 consecutive amino acids were cut into a sequence.Sequences of insufficient length are padded with X, and partial overlap of sequences is allowed. For PTM sites, the positive site should be placed in the middle of the sequence. If there is a peptide chain of less than 33 amino acids, its ends are padded to the target length before training. The validation set is randomly selected at training time according to 10% of the training set.
Key parameters for reproducibility:
  • Sequence processing: Fixed length 33 amino acids; ’X’ padding for shorter sequences
  • Fragment generation: 16 amino acids sliding step (50% overlap)
  • Dataset splitting: Training (80%), Test (20%), Validation (10% of training); Random seed: 42
For bioactive peptides (BPs), the entire sequence is processed as a single unit, while for PTMs, sequences are centered on the modification site with equal padding on both sides.
These parameters were optimized based on extensive experimentation and previous studies showing that a 33-amino acid window provides sufficient context for both local and global feature extraction while maintaining computational efficiency.
@@ 2.2.2 Embedding strategy
We employ ESM-2 (650M parameters) as our foundation model, combining it with a custom BaseEmbedding module to create hybrid protein sequence representations. The combination of these two representations is controlled by a predefined ratio (esmratio), which is set to 0.9 in this case. The final combined sequence representation is computed using the following weighted sum:
@@
$\text{R}_{\text{combined}}$ represents The combined sequence feature representation; $\alpha$ (set to 0.9) is the ’esm_ratio’, which defines the contribution of the ESM-2 embeddings; $\text{R}_{\text{ESM-2}}$ is the embedding generated by the ESM-2 model; $\text{R}_{\text{Base}}$ is the embedding generated by the BaseEmbedding module.
@@
Protein sequences are passed through the pre-trained ESM-2 model, which generates token-wise embeddings from its final layer. ESM-2 650M, based on the Transformer architecture, contains 33 layers with a hidden dimension of 1280. The model’s input layer embeds protein sequences into 1280-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence. These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function identification. The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context. The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein identification, function annotation, and structure identification.preexplanation
BaseEmbedding module
The BaseEmbedding module incorporates a task-adaptive embedding architecture that maps amino acids to a 128-dimensional representation space, subsequently transformed to 1280 dimensions to align with ESM-2 embeddings. This design enables learning of task-specific patterns for PTM sites and bioactive peptides. For comprehensive details on the architecture of the BaseEmbedding module, please refer to the Supplementary Material, Section “BaseEmbedding Module Architecture.”
@@
The embeddings generated by the ESM-2 model and the BaseEmbedding module are combined using the predefined ratio (esm_ratio = 0.9). This hybrid representation allows the model to retain the rich, pre-trained features from ESM-2 while incorporating learnable, task-specific features from BaseEmbedding. The BaseEmbedding component specifically addresses the limitation that general protein language models may not capture specialized patterns relevant to peptide functionality and PTM identifictaion, providing a complementary feature space that enhances discrimination capabilities for specific biological tasks. Furthermore, we adopt a residual connection mechanism between the two representations, preserving low-level feature information through identity mappings while enabling higher layers to focus on learning incremental feature differences. This design adheres to the ”gradient shortcut” principle, which mitigates the vanishing gradient problem in deep networks by ensuring efficient backpropagation of gradients during feature optimization. Such an approach has been shown to significantly enhance training stability and overall representational capacity in deep neural networksresnet1 ; resnet2 .
@@ 2.2.3 TransConv1d Network
The TransConv1d module combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data. The module utilizes multi-head self-attention with 8 heads and 4 Transformer encoder layers, incorporating residual connections and layer normalization for training stability. The complete architecture and implementation details of the TransConv1d network can be found in the Supplementary Material (Section: TransConv1d Network Architecture).
@@ 2.2.4 PosCNN Network
The PosCNN module integrates 1D convolutional operations with optional learnable positional encoding to extract local sequence features. The module employs a convolution layer with 64 output channels, kernel size 3, and ReLU activation, followed by adaptive average pooling and fully connected layers for feature mapping. Full details of the PosCNN network architecture and implementation can be found in the Supplementary Material under “PosCNN Network Architecture.”
@@ 2.2.5 Predict module
The forward process of the Predict module first processes the input data x through the TransConv1d and PosCNN modules to obtain two feature representations. The outputs are then adjusted in dimension and concatenated along the feature dimension, forming a unified representation that combines both global and local features. The concatenated features are then permuted to match the input format required by the subsequent convolutional layers. The combined output from TransConv1d and PosCNN is passed through a convolutional layer with 32 output channels, followed by adaptive max pooling to reduce the sequence length to 1, with dropout applied for regularization. Next, the processed features are passed through another convolutional layer with 64 output channels, followed by max pooling and dropout. Finally, the features are flattened and input into a fully connected layer to generate the final identification output. This sequence of operations helps refine the features, prevent overfitting, and produce robust identification results.